Find the area of the shaded region. The graph depicts the standard normal distribution of bone density scores with mean 0 and standard deviation 1 z=-0.86 z=1.21 The area of the shaded region is □ (Round to four decimal places as needed.)
Answer #1

Answer

we have to find the value of area or the probability when z is greater than -0.86 and less than 1.21

Using the formula P(z_1<z<z_2) = P(z<z_2)-P(z<z_1)

setting the values, we get

P(-0.86<z<1.21) = P(z<1.21)-P(z<-0.86) ....equation 1

using z distribution table (looking 1.0 on left most column and 0.21 on top most row and selecting the intersection of column and row)

we get  P(z<1.21) = 0.8869

Similarly,using z distribution table , we can write P(z<-0.86) = 1-P(z<0.86) = 1 - 0.8051 = 0.1949 setting the values in equation 1, we get P(-0.86<z<1.21) = P(z<1.21)-P(z<-0.86) = 0.8869-0.1949

this gives us P(-0.86<z<1.21) =0.6920

So, area of shaded region is 0.6920 (rounded to four decimals)